No Yes All
No 161 89 250
Yes 183 90 273
All 344 179 523

Rows: Dog, Columns: Dieting

What proportion of students are dog owners?
What proportions of students are dieting?
What proportion of dog owners are not dieting?
What proportion of non-dog owners are dieting?
Answer #1

1)

proportion of students are dog owners =273/523

2)

proportions of students are dieting=179/523
3)

proportion of dog owners are not dieting =183/273

4)

proportion of non-dog owners are dieting =89/250